# Search for high energy 5.5 MeV solar axions with the complete Borexino dataset

### v1

@@ 6 Conclusions
The signature of all these reactions is a $5.5$ MeV peak in the energy spectrum of the Borexino. No statistically significant indications of axion interactions were detected. New model-independent upper limits on the axion coupling constants to photons, electrons and nucleons were obtained:
@@
all at 90% c. l. Compared to previous work Bel12a , the limits are improved by a factor of 2 – 3, that corresponds to roughly an order-of-magnitude increase in sensitivity to the count rates.
From these model-independent limits, we derive constraints on the products $|g_{A\gamma}\times m_{A}|$ and $|g_{Ae}\times m_{A}|$ for the KSVZ-axion model and the DFSZ-axion model for different values of $\tan\beta$. The new Borexino results exclude large regions of axion-photon $g_{A\gamma}$, axion-electron $g_{Ae}$ coupling constants and axion masses $m_{A}$, and allow us to constrain the values of the axion masses in specific variants of the KSVZ- and DFSZ-axion models.
@@ 7 Acknowledgments
The Borexino program is made possible by funding from Istituto Nazionale di Fisica Nucleare (INFN) (Italy), National Science Foundation (NSF) (USA), Russian Science Foundation (RSF) (Grant No. 24-12-00046) and Ministry of Science and Higher Education of the Russian Federation (Project FSWU-2023-0073) (Russia), Deutsche Forschungsgemeinschaft (DFG), Cluster of Excellence PRISMA+ (Project ID 39083149), and recruitment initiative of Helmholtz-Gemeinschaft (HGF) (Germany), and Narodowe Centrum Nauki (NCN) (Grant No. UMO-2013/10/E/ST2/00180) (Poland). We acknowledge the generous hospitality and support of the Laboratori Nazionali del Gran Sasso (Italy).

### v2

@@ 6 Conclusions
The signature of all these reactions is a $5.5$ MeV peak in the energy spectrum of the Borexino detector. No statistically significant indications of axion interactions were detected. New model-independent upper limits on the product of the axion coupling constants to photons, electrons and nucleons were obtained:
@@
all at 90% c. l. Compared to previous work Bel12a , the limits are improved by a factor of (2 – 3), that corresponds to roughly an order-of-magnitude increase in sensitivity to the count rates.
From these model-independent limits, we derive constraints on the products $|g_{A\gamma}\times m_{A}|$ and $|g_{Ae}\times m_{A}|$ for the KSVZ-axion model and the DFSZ-axion model for different values of $\tan\beta^{*}$. The new Borexino results exclude large regions of axion-photon $g_{A\gamma}$, axion-electron $g_{Ae}$ coupling constants and axion masses $m_{A}$, and allow us to constrain the values of the axion masses in specific variants of the KSVZ- and DFSZ-axion models.
@@ 7 Acknowledgments
The Borexino program is made possible by funding from Istituto Nazionale di Fisica Nucleare (INFN) (Italy), National Science Foundation (NSF) (USA), Russian Science Foundation (RSF) (Grant No. 24-12-00046), Deutsche Forschungsgemeinschaft (DFG), Cluster of Excellence PRISMA+ (Project ID 39083149), and recruitment initiative of Helmholtz-Gemeinschaft (HGF) (Germany), and Narodowe Centrum Nauki (NCN) (Grant No. UMO-2013/10/E/ST2/00180) (Poland). We acknowledge the generous hospitality and support of the Laboratori Nazionali del Gran Sasso (Italy).